A 165 Lb. male is receiving a premixed dopamine drip (400 mg/250 mL) in the ICU. The order states that the drip can be titrated to a maximum dose of 20 mcg/kg.min. What is the maximum rate in milliliters per hour at which the drip can be run?
Answer #1

AVAILABLE QUANTITY OF DOPAMINE=400 mg/250 ml

one ml contains 400/250=1.6 mg of dopamine

maximum dose/kg/ min=20 mcg

maximum dose/ kg/hr=20*60=1200mcg per kg ie 1.2 mg per kg

weight of patient=165 lb ie 74.8 kg

maximum dose per hour= wight*dose/kg/hr ie 1.2*74.8=89.76 mg approx 90 mg..

iml contains 1.6 mg of dopamine.1 mgof dopamine is 1/1.6=0.625 ml

90 mg of dopamine is90*0.625=56.25 ml

so the maximum ml/hr of the drip is56.25 ml